# ImpliRet: Benchmarking the Implicit Fact Retrieval Challenge

### v2

@@ Abstract
Zeinab Sadat Taghavi*, Ali Modarressi*, Yunpu Ma, Hinrich Schütze Center for Information and Language Processing, LMU Munich Munich Center for Machine Learning (MCML) Correspondence: zeinabtaghavi@cis.lmu.de
@@ 5 Conclusion
We introduce ImpliRet, a benchmark for evaluating retrieval models when relevance depends on document-side reasoning on implicit facts. Unlike prior datasets that emphasize complex queries, ImpliRet shifts the reasoning burden to the documents. It covers three reasoning types – world knowledge, arithmetic, and temporal – and two discourse styles. Across sparse, dense, and KG-augmented retrievers, the best nDCG@10 is only 14.91. Even with GPT-o4-mini, given thirty documents including the gold, performance peaks at just 55.54%. These results highlight the difficulty of retrieving implicit facts and the need for models that can reason beyond surface cues.
@@ A.4.2 Multi-Speaker (Forum-Style)
Forum-Base-Question | Daily work verb |

### v3

@@ Abstract
Zeinab Sadat Taghavi*, Ali Modarressi*, Yunpu Ma, Hinrich Schütze Center for Information and Language Processing, LMU Munich Munich Center for Machine Learning (MCML) Correspondence: [zeinabtaghavi, amodaresi]@cis.lmu.de
@@ 2.2 Document-Query Pairs
Fluency and implicitness sanity check.
We drew a stratified random sample of 72 instances (query–document pairs; 3 reasoning categories × 2 discourse styles × 12 per cell) and manually assessed each for (i) fluency, (ii) implicit support for the queried fact, and (iii) absence of explicit leakage (i.e., a verbatim statement of the fact). In this sample, all documents were fluent and supported the queries implicitly; under our rubric, we observed no cases of explicit leakage. Further details are provided in Appendix A.5.
@@ 4 Results
Error Analysis
RAG has two stages—retrieval and generation—so we analyze errors along two axes. 1. Retrieval side (Rank-1 vs. Positive). For each query, we compare the retriever’s top-1 passage with the annotated positive. We analyze the 60 queries where Dragon+’s top-1 document differs from the positive ( reasoning categories discourse styles queries), yielding 120 passages (top-1 and positive per query). We categorize mis-rank reasons into four groups: (i) Word Overlap (top-1 has extra query surface tokens), (ii) Semantic Cue (similar overlap but extra topical/theme terms), (iii) Length (overlap/semantics comparable; shorter passage chosen), and (iv) Unknown (indistinguishable under our heuristics). Table 4 shows that Semantic Cue is most frequent in arithmetic queries, and Word Overlap in temporal and world knowledge. 2. Generation side (Oracle-RAG, vs. all). To isolate generation errors, we evaluate an oracle setting where the positive passage is guaranteed in context. We randomly select 60 queries () and evaluate two context sizes ( and , others selected randomly), yielding 120 cases. After reviewing outputs, we assign a single label to each incorrect answer: (i) Malformation (positive present, answer malformed), (ii) No-answer/unrelated, or (iii) Distraction (copied from a surface-similar distractor). Table 5 shows that No-answer/unrelated is most frequent, Distraction occurs mainly in temporal queries, and longer context reduces Malformation but raises No-answer/unrelated, suggesting that context length alone does not fix generation errors.
| Reasoning | Word overlap | Semantic cue | Length | Unknown |
| Arithmetic | 15% | 55% | 5% | 25% |
| Temporal | 55% | 5% | 35% | 5% |
| W. Know. | 50% | 5% | 30% | 15% |
| Reasoning | Malformation | No-answer/Unrelated | Distraction | |
| Arithmetic | 10 | 40% | 60% | 0% |
| all | 20% | 75% | 5% | |
| W. Know. | 10 | 45% | 55% | 0% |
| all | 15% | 85% | 0% | |
| Temporal | 10 | 0% | 35% | 65% |
| all | 0% | 40% | 60% |
@@ 5 Conclusion
We introduce ImpliRet, a benchmark for evaluating retrieval models when relevance depends on document-side reasoning on implicit facts. Unlike prior datasets that emphasize complex queries, ImpliRet shifts the reasoning burden to the documents. It covers three reasoning types – world knowledge, arithmetic, and temporal – and two discourse styles. Across sparse, dense, and KG-augmented retrievers, the best nDCG@10 is only 14.91. Even with GPT-o4-mini, given thirty documents including the positive, performance peaks at just 55.54%. These results highlight the difficulty of retrieving implicit facts and the need for models that can reason beyond surface cues.
@@ A.4.2 Multi-Speaker (Forum-Style)
A.5 Human Evaluation Details
We performed a small, stratified sanity check to complement automatic validation. For each reasoning category discourse style cell, we randomly selected two queries (12 instances per cell). For each selected query, candidate passages were ranked with the ReasonIR retriever, and six passages were drawn by uniform sampling from two rank strata: top–20 (3) and bottom–20 (3). This yielded 72 passages in total. We assessed each passage for (i) fluency, (ii) implicit support of their corresponding queried fact (entailed but not stated verbatim), and (iii) absence of explicit leakage. In this sample, all passages were fluent and passed the implicitness and non-leakage checks under our rubric.
@@ A.5 Human Evaluation Details
| Forum-Base-Question | Daily work verb | |||